OpenStudy (anonymous):

hich arrangement shows 34/5 , 6.7 , 6 and 2/3 , and 6.65 in order from least to greatest? giving medals plz help

OpenStudy (anonymous):

6.5

OpenStudy (anonymous):

ok....

OpenStudy (anonymous):

hold on sorry

OpenStudy (anonymous):

@abdela25 plz help me :(

OpenStudy (anonymous):

i try to help

OpenStudy (anonymous):

ok thx

OpenStudy (abdela25):

okay well to simplfy it make them all into decimals first ? to do that you divide each number by each other for example 2/3 would = 0.66

OpenStudy (anonymous):

k thx

OpenStudy (abdela25):

so now we have: 2/3=0.66 6.65 6.7 34/5=6.8

OpenStudy (anonymous):

thx

OpenStudy (abdela25):

smaller is 6.65 then 2/3 then 6.7 then 34/5

OpenStudy (abdela25):

np anytime